# Personalization of Large Language Models: A Survey

### v1

@@ 6 Taxonomy of Evaluation Metrics for Personalized LLMs
  • Intrinsic Evaluation (Sec. 6.1): Formally, intrinsic evaluation metrics can be defined as $\psi(\mathcal{D})=\{\mathbb{E}_{i}\mid\mathbb{E}_{i}(\hat{y},Y)\}$. Several widely used metrics fall into this category, including BLEU (Papineni et al., 2002), ROUGE-1 (Lin & Hovy, 2003), ROUGE-L (Lin & Och, 2004), METEOR (Banerjee & Lavie, 2005), BERTScore (Zhang et al., 2019), and Hits@K. BLEU is primarily used for text generation tasks, such as machine translation. ROUGE-1 and ROUGE-L belong to the ROUGE family (Lin, 2004), originally designed for summarization evaluation. ROUGE-1 measures unigram recall between the predicted and reference summaries, while ROUGE-L considers the longest common subsequence between them. METEOR, originally developed for machine translation evaluation, focuses on string alignment. BERTScore measures the similarity between contextual embeddings generated by the BERT model (Devlin et al., 2019). Hits@K computes the percentage of tested items that appear in the top $k$ returned answers. Higher scores in these metrics indicate better model performance.
@@ 6.1 Intrinsic Evaluation
LLMs are increasingly being used as evaluators to reduce the reliance on human labor. For example, MT-bench and Chatbot Arena (Zheng et al., 2023a) strong LLMs as judges to evaluate these models on open-ended questions. However, questions remain about how reliably LLMs can serve as evaluators. Chiang & Lee (2023) first defined LLM evaluation as the process of feeding LLMs the same instructions and questions given to human evaluators and obtaining answers directly from LLMs. Judge-Bench (Bavaresco et al., 2024) provides an empirical study comparing LLM evaluation scores with human judgments, concluding that LLMs are not yet ready to fully replace human judges in NLP tasks due to high variance in their evaluation performance. EvalGen (Shankar et al., 2024) incorporates human preferences by allowing users to grade LLM-generated prompts and code. This approach iteratively refines evaluation criteria based on user feedback, helping to validate the quality of generated content. In the specific case of code generation, the pass rate metric is often used for self-evaluation. In the context of personalization, to the best of our knowledge, there is still a lack of an LLM-as-a-judge framework specifically designed to systematically evaluate personalization applications. Developing such a framework could be a promising direction for future research. Additionally, human evaluation remains indispensable in this area. Methods like human preference judgments and pairwise comparisons are commonly used to assess the alignment between generated content and user-specific requirements. While LLMs can assist in the evaluation or annotation process (Li et al., 2023c), human evaluation remains essential for ensuring that personalized outputs truly meet user expectations, despite its practical cost.
@@ 6.2 Extrinsic Evaluation
In classification tasks, personalized LLMs classify input text (such as profiles or descriptions) into one of several candidate classes, which may involve binary or multiclass classification. For instance, personalized LLMs may match patients with suitable clinical trials, where the input includes patient health records and trial descriptions, and the output indicates whether there is a match (Yuan et al., 2023). In this context, metrics such as Recall, Precision, and F1 Score are used to evaluate the quality of the matches. These metrics are similarly applied in other classification tasks, including entity recognition and relation extraction (Tang et al., 2023). Additionally, metrics like accuracy and micro-F1 Score are often employed to assess the quality of classification outputs. For example, personalized LLMs may classify movies into specific categories based on their profiles (Salemi et al., 2023), with accuracy and micro-F1 used to measure classification performance. Beyond the tasks mentioned, a wide range of personalization tasks may require other task-specific metrics for evaluation.
@@
Table 5: Taxonomy of Datasets for Personalized LLMs. The datasets are categorized by the downstream task they have been applied to, including: short-text generation, long-text generation, recommendation, classification, dialogue, and question answering. Notably, benchmark datasets for short-text and long-text generation contain user-specific ground-truth text, while others primarily rely on task-specific labels in different formats. For each dataset, we also indicate whether data has been filtered to include only users with a substantial amount of prior interactions (e.g., users who have reviewed at least 100 products, sent at least 100 emails, or rated at least $k$ movies). In addition, we identify whether the dataset contains user-generated text, numerical attributes (such as ratings), or other categorical attributes (e.g., the genre of a movie a user watched). While attributes like movie descriptions are text-based, they are not considered user-generated content for the purpose of personalized LLM techniques. Evaluation metrics are abbreviated as follows: ROUGE-1 (R1), ROUGE-L (RL), METEOR (MET), Precision (P), Recall (R), Normalized Discounted Cumulative Gain (NDCG), Hit Ratio (HR), Accuracy (ACC), Mean Absolute Error (MAE), Root Mean Square Error (RMSE), Perplexity (PPL), and additional metrics for opinion distribution analysis such as Representativeness, Steerability, Consistency, and Wasserstein Distance (Rep/St/Con/Was). "N/A" in the Output column indicates that the target is non-textual, while in the Evaluation Metrics column, it signifies the absence of established metrics for that dataset in personalization tasks, although the dataset may have potential for personalized applications.
@@ 7 Taxonomy of Datasets for Personalized LLMs
We also categorize the datasets under the various tasks by whether they can be used to directly evaluate the personalized text generated via some personalized LLM approach or whether the text generated is indirectly evaluated for a downstream application where the goal is to demonstrate that when the text is leveraged, it improves performance of some other model, whether it is a recommendation model, classifier, among others.
@@ 7.2 Personalized Datasets without Ground-Truth Text
Personalized datasets that are useful for indirect evaluation of the generated text via a downstream application are by far the most common, as they do not require an actual set of ground-truth text written by individual users, and instead can leverage user attributes and other interaction data, to generate personalized text, and then this text is used to enhance another arbitrary model such as a recommendation approach. Such datasets that can be used in this fashion are shown in Table 5 in the following categories, including, recommendation, classification, dialogue, and question answering. By leveraging such evaluation strategy enables one to leverage any commonly used dataset that has been used for a variety of different personalization tasks such as recommendation, classification and so on. However, such criticism to this approach is that we can only demonstrate that using the text is useful for the downstream application, but not that the text was actually meaningful or relevant to the user. For instance, the generated text could be a string of random words, which we then obtain an embedding via sentence transformer, and then augment the existing user embedding with this embedding from the generated text, and demonstrate that by using this additional information, we can improve the performance of the downstream task.
Additionally, we highlight datasets that, while not originally designed for personalization tasks, contain rich user-specific information and user-generated text, making them suitable for downstream personalization tasks. Notable examples include the PRISM Alignment Dataset (Kirk et al., 2024), which links survey responses and demographic information from diverse participants to their interactions with various LLMs, and the Empathic Conversations dataset (Omitaomu et al., 2022), which features multi-turn conversations enriched with detailed demographic and personality profiles, pre- and post-conversation data, and turn-level annotations. These datasets are valuable for tasks such as personalized QA, dialogue generation, empathy-aware responses, and user-specific emotion modeling in language models.
@@ 8.1.2 Healthcare
LLMs have demonstrated remarkable proficiency in various healthcare-related tasks (Liu et al., 2023e; Wang et al., 2023b; Liu et al., 2023h; Yang et al., 2024b), paving the way for their potential integration into personalized health assistance. Belyaeva et al. (2023) introduce HeLM, a framework that enables LLMs to leverage individual-specific multimodal health data for personalized disease risk prediction. HeLM employs separate encoders to map non-text data modalities (such as tabular clinical features and high-dimensional lung function measures) into the LLM’s token embedding space, allowing the model to process multimodal inputs together. Abbasian et al. (2023) present openCHA, an open-source LLM-powered framework for conversational health agents that enables personalized healthcare responses by integrating external data sources, knowledge bases, and analytical tools. The framework features an orchestrator for planning and executing information-gathering actions and incorporates multimodal and multilingual capabilities. Building on openCHA, Abbasian et al. (2024) integrate specific diabetes-related knowledge to enhance performance in downstream tasks within the domain. Zhang et al. (2024a) introduce MaLP, a novel framework for personalizing LLMs as medical assistants. The approach combines a dual-process enhanced memory (DPeM) mechanism, inspired by human memory processes, with PEFT to improve LLMs’ ability to provide personalized responses while maintaining low resource consumption. Jin et al. (2024b) propose a Health-LLM-based pipeline with RAG to provide personalized disease prediction and health recommendations. The system extracts features from patient health reports using in-context learning, assigns scores to these features using medical knowledge, and then employs XGBoost for final disease prediction.
@@ 8.1.3 Other Domains
Beyond the two domains where personalized LLMs have been widely employed, this section explores areas with less focus but significant potential for applying personalized LLMs. In these domains, specialized LLMs or agentic frameworks are emerging, but they often lack emphasis on personalization—an aspect that could greatly enhance user experiences.
@@
As LLMs continue to advance, especially those fine-tuned on code-specific datasets (Roziere et al., 2023; Chen et al., 2021), their ability to generate high-quality code has seen significant improvement. This has spurred the development of an increasing number of AI-powered assistants aimed at enhancing the coding experience for programmers (Zhang et al., 2024b; Wang et al., 2024d; c; Xia et al., 2024). However, these applications often overlook the crucial aspect of personalization. For personalized code generation, Dai et al. (2024) propose MPCODER a novel approach for generating personalized code for multiple users that aligns with their individual coding styles. This method uses explicit coding style residual learning to capture syntax standards and implicit style learning to capture semantic conventions, along with a multi-user style adapter to differentiate between users through contrastive learning. The authors introduce a new evaluation metric called Coding Style Score to quantitatively assess coding style similarities. Personalization in coding assistance can be realized in several ways. Firstly, programmers and teams often have unique coding styles. For example, a middle-school student might prefer code that is easy to understand and well-commented, while a software engineer in a tech company may prioritize performance, scalability, and strict adherence to industry standards. A personalized LLM can learn from a user’s behavior over time and adapt their suggestions to match the user’s skill level, preferred frameworks, and commonly used coding patterns would significantly enhance their utility. Secondly, context-aware debugging is another area where personalization can make a substantial impact. Personalized LLMs could offer tailored debugging assistance based on a programmer’s typical errors and preferred debugging strategies. Thirdly, enforcing code guidelines that align with a team’s standards such as naming conventions, architectural patterns, and tool integrations is essential in collaborative environments. This ensures consistency and maintainability across the codebase, which is particularly critical in professional settings. Finally, personalized LLMs could greatly improve collaboration and code review processes by offering suggestions that consider both individual and team preferences. Achieving such levels of personalization requires advanced techniques like RAG or fine-tuning on user-specific data, enabling the model to adapt to the distinct needs and preferences of different programmers. This represents a promising direction for future research and development in AI-powered coding assistants.
@@ 8.2 Recommendation
Personalized LLMs have been extensively applied across various recommendation tasks, including direct recommendations, sequential recommendations, conversational recommendations, explainable recommendations, rating predictions, and ranking predictions (Dai et al., 2023; Du et al., 2024; Hou et al., 2024; Liu et al., 2023a; 2024c; Ji et al., 2024b). These applications aim to enhance user experiences in recommendation domains like e-commerce (Tan & Jiang, 2023; Chen et al., 2024c). According to Wu et al. (2023a), the integration of LLMs into recommendation systems can be categorized into three main approaches: (1) augmenting traditional recommendation systems, such as collaborative filtering (Schafer et al., 2007; Resnick et al., 1994), with LLM embeddings; (2) enriching traditional recommendation systems by using LLM-generated outputs as features; and (3) employing LLMs directly as recommenders in downstream recommendation tasks. These three approaches can all benefit from the personalization techniques discussed in Sec. 5. To enhance traditional recommendation systems with personalized LLMs, PALR (Yang et al., 2023) leverages LLMs to generate natural language user profiles, retrieve relevant item candidates, and rank and recommend items using a fine-tuned LLM. Chat-Rec improves traditional recommender systems by using LLMs to increase interactivity and explainability. It converts user profiles and historical interactions into prompts, allowing LLMs to learn user preferences through in-context learning and generate more personalized recommendations. For direct use of personalized LLMs as recommenders, InstructRec (Zhang et al., 2023a) frames recommendation tasks as instruction-following tasks for LLMs. It designs a flexible instruction format that incorporates user preferences, intentions, and task forms, and generates a large dataset of personalized instructions to fine-tune an LLM for recommendation tasks. Similarly, GeneRec (Wang et al., 2023a) utilizes LLMs for personalized content creation based on user instructions and feedback, aiming to complement traditional retrieval-based recommendation systems. While personalized LLMs have seen widespread application in recommendation systems, demonstrating superior performance in few-shot and zero-shot settings with enhanced explainability that addresses the cold-start problem (Liu et al., 2023a; Dai et al., 2023; Hou et al., 2024), significant challenges remain, including concerns over privacy, cost, and latency in large-scale deployments, highlighting the need for continued research and innovation

### v3

@@ 6 Taxonomy of Evaluation Metrics for Personalized LLMs
Definition 20 (Intrinsic Evaluation).
$\mathbb{E}_{i}(\hat{y},Y)=\psi(\hat{y},Y)$
Note that the ground-truth data $Y$ can represent user-written text when available or, alternatively, user preferences and reward models that capture user judgments.
Definition 21 (Extrinsic Evaluation).
$\mathbb{E}_{e}(\hat{r},r)=\psi_{a}(\hat{r},r)$
@@
  • Intrinsic Evaluation (Sec. 6.1): Formally, intrinsic evaluation metrics can be defined as $\psi(\mathcal{D})=\{\mathbb{E}_{i}\mid\mathbb{E}_{i}(\hat{y},Y)\}$. Most research on personalized text generation focuses on scenarios where ground-truth user-written text is available (Salemi et al., 2023; Kumar et al., 2024). In this setting, common evaluation metrics include BLEU (Papineni et al., 2002), ROUGE-1 (Lin & Hovy, 2003), ROUGE-L (Lin & Och, 2004), METEOR (Banerjee & Lavie, 2005), BERTScore (Zhang et al., 2019), and Hits@K. BLEU is primarily used for text generation tasks, such as machine translation. ROUGE-1 and ROUGE-L belong to the ROUGE family (Lin, 2004), originally designed for summarization evaluation. ROUGE-1 measures unigram recall between the predicted and reference summaries, while ROUGE-L considers the longest common subsequence between them. METEOR, originally developed for machine translation evaluation, focuses on string alignment. BERTScore measures the similarity between contextual embeddings generated by the BERT model (Devlin et al., 2019). Hits@k measures the percentage of test cases where the correct answer appears in the top $k$ predictions. For example, with the persona “I love hiking” and the correct response “I hike every weekend” ranked first, it counts as a hit for Hits@1, Hits@3, and Hits@5. Higher scores in these metrics indicate better model performance. However, there are scenarios where ground-truth text is not available. In such cases, model alignment with human pairwise preferences or reward models that capture user intent can serve as alternatives. However, this aspect has not been systematically explored yet.
@@ 6.1 Intrinsic Evaluation
LLMs are increasingly being used as evaluators to reduce the reliance on human labor. For example, MT-bench and Chatbot Arena (Zheng et al., 2023a) use strong LLMs as judges to evaluate these models on open-ended questions. However, questions remain about how reliably LLMs can serve as evaluators. Chiang & Lee (2023) first defined LLM evaluation as the process of feeding LLMs the same instructions and questions given to human evaluators and obtaining answers directly from LLMs. Judge-Bench (Bavaresco et al., 2024) provides an empirical study comparing LLM evaluation scores with human judgments, concluding that LLMs are not yet ready to fully replace human judges in NLP tasks due to high variance in their evaluation performance. EvalGen (Shankar et al., 2024) incorporates human preferences by allowing users to grade LLM-generated prompts and code. This approach iteratively refines evaluation criteria based on user feedback, helping to validate the quality of generated content. In the context of personalization, to the best of our knowledge, there is still a lack of an LLM-as-a-judge framework specifically designed to systematically evaluate personalization applications. Developing such a framework or constructing robust reward models that accurately capture diverse user preferences could be a promising direction for future research. Additionally, human evaluation remains indispensable in this area. Methods like human preference judgments and pairwise comparisons are commonly used to assess the alignment between generated content and user-specific requirements. While LLMs can assist in the evaluation or annotation process (Li et al., 2023d), human evaluation remains essential for ensuring that personalized outputs truly meet user expectations, despite its practical cost.
@@ 6.2 Extrinsic Evaluation
In classification tasks, personalized LLMs classify input text (such as profiles or descriptions) into one of several candidate classes, which may involve binary or multiclass classification. For instance, personalized LLMs may match patients with suitable clinical trials, where the input includes patient health records and trial descriptions, and the output indicates whether there is a match (Yuan et al., 2023). In this context, metrics such as Recall, Precision, and F1 Score are used to evaluate the quality of the matches. These metrics are similarly applied in other classification tasks, including entity recognition and relation extraction (Tang et al., 2023). Additionally, metrics like accuracy and micro-F1 Score are often employed to assess the quality of classification outputs. For example, personalized LLMs may classify movies into specific categories based on their profiles (Salemi et al., 2023), with accuracy and micro-F1 used to measure classification performance. Beyond the tasks mentioned, a wide range of personalization tasks may require other task-specific metrics for evaluation. While most of these metrics often do not directly assess personalization, improvements often implicitly reflect enhanced personalization capabilities through better utilization of user information.
@@
Table 5: Taxonomy of Datasets for Personalized LLMs. The datasets are categorized by the downstream task they have been applied to, including: short-text generation, long-text generation, recommendation, classification, dialogue, and question answering.. Notably, benchmark datasets for short-text and long-text generation contain user-specific ground-truth text, while others primarily rely on task-specific labels in different formats. For each dataset, we also indicate whether data has been filtered to include only users with a substantial amount of prior interactions (e.g., users who have reviewed at least 100 products, sent at least 100 emails, or rated at least $k$ movies). In addition, we identify whether the dataset contains user-generated text, numerical attributes (such as ratings), or other categorical attributes (e.g., the genre of a movie a user watched). While attributes like movie descriptions are text-based, they are not considered user-generated content for the purpose of personalized LLM techniques. Evaluation metrics are abbreviated as follows: ROUGE-1 (R1), ROUGE-L (RL), METEOR (MET), Precision (P), Recall (R), Normalized Discounted Cumulative Gain (NDCG), Hit Ratio (HR), Accuracy (ACC), Mean Absolute Error (MAE), Root Mean Square Error (RMSE), Perplexity (PPL), and additional metrics for opinion distribution analysis such as Representativeness, Steerability, Consistency, and Wasserstein Distance (Rep/St/Con/Was). "N/A" in the Output column indicates that the target is non-textual, while in the Evaluation Metrics column, it signifies the absence of established metrics for that dataset in personalization tasks, although the dataset may have potential for personalized applications.
@@ 7 Taxonomy of Datasets for Personalized LLMs
We also categorize the datasets based on whether they are used for direct evaluation of personalized text generation through reward signals that reflect user preferences or for indirect evaluation in a downstream application. In the former case, as most existing datasets for personalized text generation rely on ground-truth text for evaluation, our discussion primarily focuses on these scenarios. In the latter case, the goal is to demonstrate that incorporating the generated text improves the performance of a downstream task, such as a recommendation model, classifier, or other functional systems.
@@ 7.2 Personalized Datasets without Ground-Truth Text
Personalized datasets that are useful for indirect evaluation of the generated text via a downstream application are by far the most common, as they do not require an actual set of ground-truth text written by individual users, and instead can leverage user attributes and other interaction data, to generate personalized text, and then this text is used to enhance another arbitrary model such as a recommendation approach. Such datasets that can be used in this fashion are shown in Table 5 in the following categories, including, recommendation, classification, dialogue, and question answering. Leveraging such category of evaluation strategy enables one to leverage any commonly used dataset that has been used for a variety of different personalization tasks such as recommendation, classification, and so on. However, a criticism of this approach is that we can only demonstrate that using the text is useful for the downstream application, but not that the text was actually meaningful or relevant to the user. On the other hand, for instance, generated intermediate tokens or embeddings can be used to augment user embeddings. The augmented embeddings are then shown to improve the performance of downstream tasks, highlighting the potential value of this approach while acknowledging its inability to validate user-specific relevance directly.
Additionally, we highlight datasets that, while not originally designed for personalization tasks, contain rich user-specific information and user-generated text, making them suitable for downstream personalization tasks. Notable examples include the PRISM Alignment Dataset (Kirk et al., 2024), which links survey responses and demographic information from diverse participants to their interactions with various LLMs, and the Empathic Conversations dataset (Omitaomu et al., 2022), which features multi-turn conversations enriched with detailed demographic and personality profiles, pre- and post-conversation data, and turn-level annotations. These datasets are valuable for tasks such as personalized QA, dialogue generation, empathy-aware responses, and user-specific emotion modeling in language models. We believe that although these datasets do not primarily use user-generated text as ground truth for evaluation, their rich user-specific information has the potential to capture diverse preferences. Developing effective reward models based on this information could be valuable for evaluating personalized text generation in cases where ground-truth user-generated text is unavailable, an area that remains largely unexplored.
@@ 8.1.2 Healthcare
LLMs have demonstrated remarkable proficiency in various healthcare-related tasks (Liu et al., 2023e; Wang et al., 2023c; Liu et al., 2023h; Yang et al., 2024b), paving the way for their potential integration into personalized health assistance. Belyaeva et al. (2023) introduce HeLM, a framework that enables LLMs to leverage individual-specific multimodal health data for personalized disease risk prediction. HeLM employs separate encoders to map non-text data modalities (such as tabular clinical features and high-dimensional lung function measures) into the LLM’s token embedding space, allowing the model to process multimodal inputs together. Abbasian et al. (2023) present openCHA, an open-source LLM-powered framework for conversational health agents that enables personalized healthcare responses by integrating external data sources, knowledge bases, and analytical tools. The framework features an orchestrator for planning and executing information-gathering actions and incorporates multimodal and multilingual capabilities. Building on openCHA, Abbasian et al. (2024) integrate specific diabetes-related knowledge to enhance performance in downstream tasks within the domain. Zhang et al. (2024a) introduce MaLP, a novel framework for personalizing LLMs as medical assistants. The approach combines a dual-process enhanced memory (DPeM) mechanism, inspired by human memory processes, with PEFT to improve LLMs’ ability to provide personalized responses while maintaining low resource consumption. Jin et al. (2024b) propose a Health-LLM-based pipeline with RAG to provide personalized disease prediction and health recommendations. The system extracts features from patient health reports using in-context learning, assigns scores to these features using medical knowledge, and then employs XGBoost for final disease prediction.
@@ 8.1.3 Other Domains
Beyond the two domains where personalized LLMs have been widely employed, this section explores areas with less focus but significant potential for applying personalized LLMs. In these domains, specialized LLMs or language agent frameworks are emerging, but they often lack emphasis on personalization—an aspect that could greatly enhance user experiences.
@@
As LLMs continue to advance, especially those fine-tuned on code-specific datasets (Roziere et al., 2023; Chen et al., 2021), their ability to generate high-quality code has seen significant improvement. This has spurred the development of an increasing number of AI-powered assistants aimed at enhancing the coding experience for programmers (Zhang et al., 2024b; Wang et al., 2024e; d; Xia et al., 2024). However, these applications often overlook the crucial aspect of personalization. For personalized code generation, Dai et al. (2024) propose MPCODER a novel approach for generating personalized code for multiple users that aligns with their individual coding styles. This method uses explicit coding style residual learning to capture syntax standards and implicit style learning to capture semantic conventions, along with a multi-user style adapter to differentiate between users through contrastive learning. The authors introduce a new evaluation metric called Coding Style Score to quantitatively assess coding style similarities. Personalization in coding assistance can be realized in several ways. Firstly, programmers and teams often have unique coding styles. For example, a middle-school student might prefer code that is easy to understand and well-commented, while a software engineer in a tech company may prioritize performance, scalability, and strict adherence to industry standards. A personalized LLM can learn from a user’s behavior over time and adapt their suggestions to match the user’s skill level, preferred frameworks, and commonly used coding patterns would significantly enhance their utility. Secondly, context-aware debugging is another area where personalization can make a substantial impact. Personalized LLMs could offer tailored debugging assistance based on a programmer’s typical errors and preferred debugging strategies. Thirdly, enforcing code guidelines that align with a team’s standards such as naming conventions, architectural patterns, and tool integrations is essential in collaborative environments. This ensures consistency and maintainability across the codebase, which is particularly critical in professional settings. Finally, personalized LLMs could greatly improve collaboration and code review processes by offering suggestions that consider both individual and team preferences. Achieving such levels of personalization requires advanced techniques like RAG or fine-tuning on user-specific data, enabling the model to adapt to the distinct needs and preferences of different programmers. This represents a promising direction for future research and development in AI-powered coding assistants.
@@ 8.2 Recommendation
Personalized LLMs have been extensively applied across various recommendation tasks, including direct recommendations, sequential recommendations, conversational recommendations, explainable recommendations, rating predictions, and ranking predictions (Dai et al., 2023; Du et al., 2024; Hou et al., 2024; Liu et al., 2023a; 2024d; Ji et al., 2024b). These applications aim to enhance user experiences in recommendation domains like e-commerce (Tan & Jiang, 2023; Chen et al., 2024c). According to Wu et al. (2023a), the integration of LLMs into recommendation systems can be categorized into three main approaches: (1) augmenting traditional recommendation systems, such as collaborative filtering (Schafer et al., 2007; Resnick et al., 1994), with LLM embeddings; (2) enriching traditional recommendation systems by using LLM-generated outputs as features; and (3) employing LLMs directly as recommenders in downstream recommendation tasks. These three categories of approaches can all benefit from the personalization techniques discussed in Sec. 5. To enhance traditional recommendation systems with personalized LLMs, PALR (Yang et al., 2023) leverages LLMs to generate natural language user profiles, retrieve relevant item candidates, and rank and recommend items using a fine-tuned LLM. Chat-Rec improves traditional recommender systems by using LLMs to increase interactivity and explainability. It converts user profiles and historical interactions into prompts, allowing LLMs to learn user preferences through in-context learning and generate more personalized recommendations. For direct use of personalized LLMs as recommenders, InstructRec (Zhang et al., 2023a) frames recommendation tasks as instruction-following tasks for LLMs. It designs a flexible instruction format that incorporates user preferences, intentions, and task forms, and generates a large dataset of personalized instructions to fine-tune an LLM for recommendation tasks. Similarly, GeneRec (Wang et al., 2023b) employs LLMs for personalized content creation based on user instructions and feedback, aiming to complement traditional retrieval-based recommendation systems. While personalized LLMs have seen widespread application in recommendation systems, demonstrating superior performance in few-shot and zero-shot settings with enhanced explainability that addresses the cold-start problem (Liu et al., 2023a; Dai et al., 2023; Hou et al., 2024), significant challenges remain, including concerns over privacy, cost, and latency in large-scale deployments, highlighting the need for continued research and innovation